### 20 Jan 2024 v1

# StickerConv: Generating Multimodal Empathetic Responses from Scratch

Abstract
Stickers, while widely recognized for enhancing empathetic communication in online interactions, remain underexplored in current empathetic dialogue research. In this paper, we introduce the Agent for StickerConv (Agent4SC), which uses collaborative agent interactions to realistically simulate human behavior with sticker usage, thereby enhancing multimodal empathetic communication. Building on this foundation, we develop a multimodal empathetic dialogue dataset, StickerConv, which includes 12.9K dialogue sessions, 5.8K unique stickers, and 2K diverse conversational scenarios, specifically designs to augment the generation of empathetic responses in a multimodal context. To leverage the richness of this dataset, we propose PErceive and Generate Stickers (PEGS), a multimodal empathetic response generation model, complemented by a comprehensive set of empathy evaluation metrics based on LLM. Our experiments demonstrate PEGS’s effectiveness in generating contextually relevant and emotionally resonant multimodal empathetic responses, contributing to the advancement of more nuanced and engaging empathetic dialogue systems. Our project page is available at https://neu-datamining.github.io/StickerConv.
StickerConv: Generating Multimodal Empathetic Responses from Scratch
[…]
footnotetext: Corresponding author.**footnotetext: Equal contribution.
1 Introduction
Increasing research indicates that utilizing stickers in online chats can effectively alleviate stress, augment personal happiness, and notably boost empathy Kariko and Anasih (2019); Akram and Drabble (2022); Jiang et al. (2023). Prior studies on stickers primarily concentrated on sentiment analysis Ge et al. (2022) and recommendation systems Gao et al. (2020); Zhang et al. (2022), overlooking their vast potential in empathetic response generation. Most empathetic response generation tasks focus solely on textual modality Fu et al. (2023); Lee et al. (2022); Jiang et al. (2023), yet stickers in chats convey more abundant and intuitive emotional information, enhancing the expressiveness and emotional depth of responses.
Figure 1: An example of multimodal conversation in the StickerConv. Both parties can utilize the stickers to express their emotions, which enhances interactivity and expression. Assistant can empathize with the user according to the conversation (green text).
Acknowledging stickers as a crucial non-textual modality and incorporating them into empathetic response generation offers novel perspectives and methodologies for empathetic dialogue tasks Zech and Rimé (2005); Rashkin et al. (2019). Specifically, integrating text and stickers paves the way for a new multimodal empathetic communication approach, targeting the generation of more diverse and high-quality empathetic responses.
A primary challenge in integrating stickers into empathetic response generation is developing a high-quality dataset to support this innovative multimodal communication. Traditional dataset construction methods, including expensive human crowdsourcing, privacy issues Xu et al. (2022), the difficulty in acquiring data from social software companies, and the variable quality of Internet-crawled data Foerderer (2023), face significant limitations.
To address this, leveraging large language models (LLMs) for dataset construction is proposed. LLMs, with their extensive world knowledge and text processing capabilities (OpenAI, 2023; Touvron et al., 2023a, b), demonstrate near-human annotation abilities Wang et al. (2023d); Pangakis et al. (2023). However, LLMs have limitations in empathetic tasks, excelling in responding to explicit human instructions but lacking in proactivity Wu et al. (2023a); Yin et al. (2023), a critical aspect of empathy. Empathy necessitates understanding others’ emotions and the ability to actively express support and understanding Zech and Rimé (2005); Sharma et al. (2020a). To mitigate this, we introduce a multi-agent system based on LLMs, Agent for StickerConv (Agent4SC). This system, through inter-agent interactions, utilizes stickers to simulate human-like dialogue scenarios. It not only generates text responses but also strategically selects suitable stickers, thereby effectively enhancing empathy.
Based on the superior capabilities and interactions of Agent4SC, we built a multimodal empathetic dataset, StickerConv, that comprise 12.9K dialogue sessions and 5.8K unique stickers. StickerConv boasts an average of 5.22 stickers per dialogue session. Figure 1 depicts an example of conversations in our dataset. To the best of out knowledge, this is the first multimodal empathetic dialogue dataset, with the particular utility of sticker as non-textual modal information to better facilitate empathy.
Although Agent4SC exhibits sound multimodal empathetic dialog response generation capability, it is limited by expensive inference costs, while the retrieval approach leads to poor performance in new scenarios. To further advance the research on multimodal empathetic dialog, based on StickerConv, we develop a end-to-end large multimodal dialog model, PEGS, with the ability to PErceive and Generate Stickers, and has a significantly lower deployment and inference cost. Beyond the general ability to generate textual empathetic responses, PEGS receives multimodal inputs and autonomously generates stickers based on the emotional and contextual aspects of the dialogue at the appropriate moment. Furthermore, to simulate human communications on social media in the real world, our model supports interleaved multiple image and text inputs.
Due to potential misalignments between the modality quantities of predicted responses and those of the golden responses, reliance on a singular modal metric may lead to significant distortions in evaluation results. Empathy, being inherently subjective, presents challenges in its direct quantification. In this context, we introduce a novel multimodal empathetic response evaluation method. This method enables the integration of multimodal information, emphasizing three critical dimensions: empathy, consistency, and preference ranking. Leveraging the world knowledge and anthropomorphic capabilities of LLMs, our method offers robust support for evaluating multimodal empathetic responses.
In conclusion, the main contributions of this work are as follows:
  • We introduce Agent for StickerConv (Agent4SC), a LLM-based multi-agent system, integrates stickers intelligently into empatheric dialogues. Agent4SC ensures the stickers are contextually consistent, varied, and empathetically aligned with human interaction. Leveraging the Agent4SC, we built a multimodal empatheric dialogue dataset, StickerConv, comprising 12.9K dialogue sessions and 5.8K unique stickers.
  • We design PErceive and Generate Stickers (PEGS), an empathetic dialogue response generation model that, within its end-to-end framework, intuitively incorporates stickers based on the emotional and contextual dynamics of a dialogue. PEGS adeptly processes multimodal inputs, generating empathetic textual responses and appropriately using stickers to enrich these replies.
  • We propose a multimodal empathetic response evaluation approach. This method integrates multimodal information and concentrates on three principal dimensions: empathy, consistency, and preference ranking.
2 Related Work
2.1 Empathetic Response Generation
The purpose of empathetic response generation is to enable machines to understand and respond to human emotions and feelings. In this field, The creation of the EMPATHETICDIALOGUES dataset Rashkin et al. (2019) laid the foundation. Subsequently, the empathetic listener mixture model Lin et al. (2019), which emphasizes understanding and appropriately responding to different emotions, and studies focusing on emotion imitation and emotion cause words to enhance the effectiveness of empathetic dialogue Majumder et al. (2020) have achieved notable results. Additionally, the introduction of the EPITOME conceptual framework Sharma et al. (2020b), a classification system for empathetic response intents Welivita and Pu (2020), research on generating empathetic responses based on commensense knowledge Sabour et al. (2021), and the proposal of the multifactorial hierarchical framework CoMAE Zheng et al. (2021) have provided new perspectives and methods for the field.
Despite these advancements, the research in this domain is not without its limitations. Primarily, the field has been domainated by small-scale models, with limited application of large language models, such as ChatGPT OpenAI (2023), in the generation of empathetic responses Lee et al. (2022); Zhou et al. (2023b). Secondly, the focus has predominantly been on text-based empathetic responses, thus neglecting the rich potential of multimodal information, such as image. Emerging studies underscore the significance of such multimodal data, especially stickers, in alleviating stress, enhancing well-being, and fostering emotional resonance and social connectivity Kariko and Anasih (2019); Akram and Drabble (2022); Jiang et al. (2023). Finally, the subjectivity and multidimensionality of empathetic tasks make it challenging to quantitatively assess their effectiveness Fu et al. (2023); Lee et al. (2022), especially in the case of multimodal empathetic responses, posing a significant challenge in the field of empathetic response generation.
2.2 Large Multimodal Models
Multimodal Perception
Large language models (LLMs), such as ChatGPT (OpenAI, 2023), LLaMA (Touvron et al., 2023a, b), demonstrate powerful capabilities in dialog interaction and instruction following, and recent researches have extended LLMs to multimodal domains. Flamingo (Alayrac et al., 2022) exhibits promising zero-shot and few-shot multimodal understanding capability by adding a cross-attention layer to connect the frozen vision encoder with the LLM. BLIP (Li et al., 2023b; Dai et al., 2023), MiniGPT-4 (Zhu et al., 2023; Chen et al., 2023) and LLaVA (Liu et al., 2023b, a) bridge the frozen vision encoder and the LLM through a small intermediate model.
Multimodal Generation
The birth and development of the Diffusion Models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Rombach et al., 2022) significantly contribute to the evolution of image generation. Koh et al. (2023) for the first time explore the mapping of the output of LLMs into the input space of the vision decoder, empowering LLMs with image generation capability. GILL (Koh et al., 2023), NExT-GPT (Wu et al., 2023b), MiniGPT-5 (Zheng et al., 2023) align LLMs to frozen vision decoders through an encoder-decoder transformer, while SEED (Ge et al., 2023) employs a learnable Q-Former. In contrast to previous works which favor realistic images, our target is to generate stickers, which are abstract and exhibit distinct emotional tendencies.
2.3 LLM-Based Agents
In the realm of artificial intelligence, Large Language Model (LLM)-based agents epitomize a cutting-edge paradigm, leveraging the robust computational prowess of LLMs. These agents are not merely confined to problem-solving and task execution; their capabilities extend to intricate reasoning and sophisticated interactions, as elucidated in recent studies Wang et al. (2023b); Li et al. (2023a). The versatility of LLM agents is further demonstrated through their adaptability in diverse domains such as software development Qian et al. (2023) and scientific research Boiko et al. (2023).
A remarkable aspect of LLM-based agents is their ability to emulate complex human social behaviors. This includes a spectrum of interactions ranging from social engagements Park et al. (2023); Tu et al. (2023), to simulating multifaceted scenarios like elections Argyle et al. (2022), debates Wang et al. (2023a), and consumer activities Wang et al. (2023c), Such capabilities underscore the potential of LLM agents in understanding and replicating human social dynamics.
The most transformative application of LLM-based agents, however, lies in their utility for generating training data for diverse AI models. Studies Wang et al. (2023d); Peng et al. (2023); Zhou et al. (2023a) have been instrumental in showcasing the efficacy of these agents in producing high-quality data, paralleling human annotators, yet at a substantially lower cost. This breakthrough has redefined the paradigm for creating training datasets, offering a more practical and cost-effective solution in the field of AI model development.
Figure 2: The Overview of Agent for StickerConv.
3 Agent for StickerConv
The limitations of Large Language Models (LLMs) and Language Model Multimodalities (LMMs), are twofold: (1) They struggle with capturing nuanced human emotions and providing empathetic responses. (2) While adept at responding to explicit human instructions, they lack the initiative to act beyond these directives.
This inherent disadvantage hinders their performance in tasks that require empathy, making them less effective in this area compared to others. To bridge this gap, we propose the Agent for StickerConv (Agent4SC), a novel LLM-based multi-agent. This system is designed to simulate human dialogue by leveraging the collaborative efforts of multiple modules, along with the use of stickers, to produce emotional and diverse empathetic responses.
The architecture of Agent4SC is depicted in Figure 2 and comprises two key agents. The User Agent, forming the core of the system, is an intricate assembly of several modules. It replicates the human processes of observation and reflection through the interplay between its Memory and Plan modules. This design effectively addresses the LLMs’ shortcomings in comprehending subtle human emotions. The insights gained from observation and reflection are articulated via the Action module. Consequently, the Agent is more adept at generating human-like emotional responses. The entire process is orchestrated by the Profile module, which imparts distinctiveness in reflection and behavior to different User Agents.
Furthermore, Agent4SC adopts a pioneering approach by treating stickers as a Tool of communication. Agents autonomously select appropriate stickers for use, mirroring human behavior in conversations. This function is facilitated through the integrated process of observation, reflection, and action. The User Agent ensures consistent performance and quality control with the support of the Manager Agent. Agent4SC is not just a static tool but a dynamic participant in conversations, capable of grasping the subtleties of human emotions and the appropriate moments for empathetic interaction.
3.1 Profile Module
The profile module serves as the foundation of the Agent4SC, shaping the personality and behavioral patterns of users in empathetic interactions. It comprises two components: Persona and Situation. The Persona section vividly portrays the user’s character traits, life background, and experiences, whereas the Situation part focuses on the user’s current circumstances, emotional state, and potential challenges they may encounter.
To develop rich and diversified user profiles, we initially manually crafted profiles with a variety of emotioinal distributions, subsequently expanding them to 2K unique profiles through the SELF-INSTRUCT Wang et al. (2023d) method by ChatGPT-3.5-turbo. Our constructed profile archive features a comprehensive emotional distribution, as illustrated in Figure 5. We believe that emotionally rich and content-diverse user profiles will enable the agent system to more authentically simulate human responses in empathetic dialogues.
3.2 Tool Module
To effectively simulate human behavior in using stickers for communication, we developed a Tool Module, converting the SER30K dataset into a functional tool. This dataset comprises 1,887 themes and 30,739 emotion-labeled stickers (Liu et al., 2022), such as happiness, anger, neutral, disgust, fear, sadness and surprise. In integrating it with the agent system, we encountered several challenges: (1) The dataset contains a significant number of stickers unsuitable for casual conversation or empathetic communication. (2) Stickers are annotated with a single emotional tag from a set of seven, lacking comprehensive content descriptions and a detailed emotional analysis. This limitation hinders the agent’s ability to appropriately utilize stickers based solely on these tags. (3) Humans selecting stickers in conversation typically do not choose from a database containing tens of thousands of sticker. Instead, they prefer to pick from a customized and smaller-scale sticker library.
To overcome these challenges, we design a series of sophisticated sticker processes to refashion sticker into an effective tool for accurate agent application:
Filter
Leveraging the LLaVa-v1.5-13b (Liu et al., 2023a), we crafted a nuanced three-step Chain of Thought Wei et al. (2023) procedure to meticulously scrutinize each sticker in SER30K, as shown in Figure 9. This approach entailed the model initially summarizing the sticker’s content, subsequently assessing its relevance to casual conversation, and finally determining suitability through iterative responses. This strategy not only affirmed selection accuracy but also mitigated model-induced randomness and hallucination discrepancies.
Annotation
For filtered stickers, a second iteration of the CoT process was executed using the same model. This phase involved responding to five pivotal queries, covering aspects like the sticker’s content, emotional conveyance, applicable scenarios, emotional tone, and potential satirical or humorous connotations (shown in Figure 9). These responses yielded comprehensive insights into each sticker, subsequently formatted into Q&A pairs, awaiting further processing.
Knowledge
Post-acquisition of intricate sticker details, the ChatGPT-3.5-Turbo model was engaged for an additional processing phase. Relying on the established Q&A pairs, this model adeptly deduced each sticker’s specific contextual suitability, culminating in the distillation of three fundamental knowledge: description, emotion, recommendation. These synthesized insights are instrumental in enriching agent’s understanding of sticker, and enable them to apply stickers correctly.
Split
In this phase of refining the SER30K sticker dataset for integration into our agent system, we focus on dividing the database into smaller, more manageable subsets. Each subset is transformed into a vector database, where the embeddings represent a unique blend of the original emotional labels of the stickers and the enriched emotional knowledge extracted in the previous steps.
After the above-mentioned processes, the original SER30K dataset has been transformed into a series of refined, smaller-scale vector databases. However, a crucial step is still required to make it a directly usable tool for agents. To address this, we have defined a specific module operation, termed "Retrieve". In this process, when an agent intends to use a sticker, it queries its accessible sub-vector database with its intended use and desired emotional expression of the sticker as search criteria. The system then retrieves and returns the top-K stickers with the highest similarity.
3.3 Plan Module
The Plan module is a core component of the Agent4SC, responsible for orchestrating the behavior of agents in empathetic interactions. To ensure that agents mimic human behavior in using stickers, we have specially designed three key actions to ensure the timely use of stickers, their coherence with the context, and their effective role in empathy.
  • Intention: Under the control of Profile module, the agent assesses the potential for using emotive stickers based on interaction history and observational data. The prompt template for this action is shown in Figure 11.
  • Query: If the agent intends to use a sticker, it will describe the emotional content of the sticker in detail. This description will be influenced by the Profile module, observation, and the specific intent behind using the sticker. The prompt template for this action is shown in Figure 12.
  • Select: The agent sends the query generated in the previous step to the sticker database accessible to it, retrieving the top-K stickers with the highest similarity. Subsequently, the agent combines the candidated sticker’s index, description, emotion and recommendation into a list. From this list, the agent selects the sticker it wishes to use. The prompt template for this action is shown in Figure 13.
3.4 Memory Module
In empathetic interactions, the coherence of communication and the consistency of personality traits are fundamental to the quality of responses. This necessitates that agents not only learn and provide feedback through observation during the interaction process, but also maintain long-term consistency in their personality traits. We adopted a strategy similar to that of Generative Agent Park et al. (2023): for short-term memory, we update the agent’s core traits based on the profile and observations; for long-term memory, the agent combines continuous interaction with external entities and regular self-reflection to maintain consistency.
3.5 Action Module
The Action module serves as the window and bridge for communication with the external world. It is responsible for receiving instructions and inputs from other parts of the system and executes two main actions based on this information:
  • Message: generates text responses based on the Agent’s Profile and Memory module.
  • Sticker: engages in sticker usage as dictated by the Plan module.
3.6 Manager Agent
The Manager Agent, a critical componet in enhancing the performance of the Agent4SC, plays a significant role in ensuring interaction quality and dialogue consistency. This agent comprises two main modules:
Quality Assurance
This module is crucial for upholding high dialogue standards, It accomplishes this by: (1) Evaluating the relevance of stickers within the conversation, ensuring that they contribute meaningfully to the dialogue. (2) Regulating the use of stickers to prevent their overuse, thus maintaining the integrity and flow of the conversation. (3) Performing regular checks on dialogue quality, with a special focus on the appropriateness of response length and the avoidance of redundant information.
Process Control
This module takes charge of guiding the dialogue’s progression. It leverages historical interactions as a reference to make informed decisions on the optimal timing for concluding dialogues.
Figure 3: The chart of emotional distribution in the choice of stickers between users and the system. This chart revealed a striking trend: users have a significant preference for stickers that convey negative emotions, in contrast to the system. The system predominantly utilizes stickers expressing neutral and positive emotions. This comparison not only reflects the distinct emotional expression preferences between users and the system but also highlights the system’s active and supportive role in interactions.
3.7 The StickerConv Dataset
The Agent4SC facilitates empathetic interaction with humans through the collaborative interplay of its modules, and possesses the capability to utilize stickers. Leveraging this, we propose the construction of a large-scale multimodal empathetic dialogue dataset, StickerConv, with stickers serving as a non-textual modality.
Figure 4: The overview architecture of PEGS. The input images are jointly encoded by image encoder, Q-Former and a linear layer. Vicuna is used as the language model. To empower the language model to generate images, the output of the LLM is first mapped into the input space of the image decoder through a feature mapper, and then the frozen image decoder is employed to generate images.
In the dataset creation process, two roles are designated: User and System. The User, acting as the dialogue initiator, is assigned a profile randomly selected from 2,000 personality profiles generated as described in Section 3.1. The System, in the roles of a listener and empathizer, is assigned a predefined positive profile. Regarding the sticker tool, we have processed and subdivided the sticker dataset into 100 smaller vector databases. In each dialogue session, both User and System have access to one randomly assigned vector database from this pool. This arrangement is designed to better simulate human sticker usage behavior while enhancing the diversity and richness of stickers. It is important to note that vector databases numbered 1 to 80 are exclusively employed for the creation of training and cross-validation sets, whereas those numbered 81 to 100 are reserved for the construction of the test set.
StickerConv consists of 12,931 dialogue sessions, 67,505 stickers (unique 5.8K stickers) and over 2K conversation topic. The detailed statistical information of our dataset is shown in Table 4. On average, each dialogue session in StickerConv has 5.22 stickers and 5.49 turns. The emotional label distribution of the stickers used by User and System is shown in Figure 3. This analysis reveals notable differences in the emotional distribution of sticker usage, attributable to the distinct roles of the User and the System.
4 PEGS
We devise an empathetic large multimodal dialog model, PEGS, with the ability to PErceive and Generate Stickers. Figure 4 illustrates the overall structure of our model. Technically, we utilize ViT-G/14 from EVA-CLIP (Fang et al., 2023), Q-Former from BLIP-2 $\text{OPT}_{\text{2.7B}}$ (Li et al., 2023b), and a linear layer to encode images. Vicuna (Chiang et al., 2023), a widely used language model in LMMs, is employed for language modeling. We adopt Stable Diffusion (SD) (Rombach et al., 2022) as our image decoder.
4.1 Multimodal Input Perception
Multimodal Encoding
With reference to existing works (Li et al., 2023b; Dai et al., 2023; Zhu et al., 2023), we convert the multimodal inputs into feature vectors that can be solved by LLM. Specially, each text token is embedded into a vector $e_{t}\in\mathbb{R}^{1\times d}$, while each image is first encoded by a pre-trained vision encoder, and then a aligned feature vector $e_{v}\in\mathbb{R}^{32\times d}$ is obtained via Q-Former and a linear projection layer.
Training
Our effort is built on BLIP-2 (Li et al., 2023b), which has captured vision-language knowledge from numerous aligned image-text pairs. Zhu et al. (2023) have proven the effectiveness that freezing the pre-trained vision encoder, Q-Former, and LLM, while only pre-training a linear projection layer. As shown in Figure 4(a), we train the input linear projection layer using LAION115M, a sub-dataset of LAION400M (Schuhmann et al., 2021). The pre-training process undergoes 80,000 steps with a batch size of 32.
4.2 Multimodal Output Generation
Expanding Vocabulary
Given original LLM vocabulary $V$ of size $|V|$, to enable the LLM to determine the timing of generating visual responses and to obtain feature vectors for image generation, we extend the vocabulary with an additional tokens set $V_{\mathrm{img}}=\{\mathrm{[IMG1]},\mathrm{[IMG2]},\ldots,\mathrm{[IMG\{}r\mathrm{\}]}\}$. Therefore, the extended vocabulary $V^{*}=V\cup V_{\mathrm{img}}$. We denote the original word embedding matrix as $E\in\mathbb{R}^{|V|\times d}$. For the embeddings matrix $E^{*}\in\mathbb{R}^{|V^{*}|\times d}$ of the extended vocabulary $V^{*}$, the embeddings $E_{\mathrm{img}}\in\mathbb{R}^{r\times d}$ of the added special token $V_{\mathrm{img}}$ are randomly initialized and the embeddings $E$ of the original token are preserved:
$\centering E^{*}[0:|V|,:]=E\@add@centering$
Text Generation
Receiving multimodal inputs, the target is to generate joint sequences of text tokens and visual tokens $\{\mathrm{[IMG\{}i{\}]}\}_{i=1}^{r}$. Specifically, the generated token can be represented as $U=\{u_{1},\ldots,u_{k}\}$, where $u_{i}\in V^{*}$. The loss function $\mathcal{L}_{\mathrm{lm}}$ is defined as:
$\centering\mathcal{L}_{\mathrm{lm}}=-\sum_{i=1}^{k}\log p(u_{i}|s,u_{1},\ldots,u_{i-1};\theta,E_{\mathrm{img}})\@add@centering$
where $s=\{e_{m}^{(1)},e_{m}^{(2)},\ldots,e_{m}^{(l)}\}$ and $m\in\{t,v\}$ denoting the modality.The original LLM weights $\theta$ are kept frozen, and we only update $E_{\mathrm{img}}$.
Image Generation
Inspired by GILL (Koh et al., 2023), we align the hidden states $h_{\mathrm{gen}}$ corresponding to the output visual tokens $\{\mathrm{[IMG\{}i{\}]}\}_{i=1}^{r}$ into the input space of the image decoder. Specifically, we connect them through a feature mapper module, containing two linear layers and a 4-layer encoder-decoder transformer model with a learnable queries feature $q$. For the given image caption $c$ (and its emotion $e$ if available), our target is to minimize the MSE loss between their embeddings derived from the frozen pre-trained SD text encoder $\tau$ and the projected representations:
$\centering\mathcal{L}_{\mathrm{gen}}=\lVert\theta_{\mathrm{Mapper}}(h_{\mathrm{gen}},q)-\tau(c,[e])\rVert_{2}^{2}\@add@centering$
Training
Distinguished from realistic images, stickers have abstract nature and rich emotional expressions. Generic text-to-image models perform well in various scenarios, but they are flawed in sticker generation. Constrained by limited sticker data, there are not enough image-text pairs to turn the base text-to-image model to the style of stickers, and to pre-train for multimodal generation. As a remedy, we adopt a variant of CuteYukiMix [^1] https://civitai.com/models/28169/cuteyukimixadorable-style, a SD model trained on cute cartoon data, as the pre-trained text-to-image model. In the pre-training phase, our objective is to align the representations with the text embeddings (Figure 4(b)), so we train our model on the captions from LAION115M. The pre-training process undergoes 50,000 steps with a batch size of 32.
However, pre-trained SDs (focusing on obvious features, such as characters’ bodies, costumes) are insensitive to sentiment conditions, and it is difficult to directly perform end-to-end instruction tuning to learn emotion elements, we inserted a supervise fine-tuning (SFT) procedure on the sticker-text pairs with emotion to efficiently cause them to focus on emotion elements.
4.3 Joint Learning
Pre-trained LLMs have powerful text interaction capability, but perform poorly on open-domain conversations. We further utilize the constructed StickerConv for jointly fine-tuning of the entire model to achieve the capabilities of open-domain empathetic multimodal perception and generation. The instruction prompts we used are shown in Appendix C.4. As shown in Figure 4(c), we train our model in an end-to-end manner, using LoRA (Hu et al., 2021) to synchronize the update of a limited number of parameters in the LLM with the input linear projection layer and the feature mapper. The overall loss function $\mathcal{L}$ can be represented as:
$\centering\mathcal{L}=\lambda_{1}\times\mathcal{L}_{\mathrm{lm}}+\lambda_{2}\times\mathcal{L}_{\mathrm{gen}}\@add@centering$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. For $\mathcal{L}_{\mathrm{gen}}$, one of its inputs — emotion $e$ is available.
5 Experiments
Experiments are conducted on StickerConv. We perform response predictions for all turns of each dialogue and consider all previous turns as context.
Models IP EX ER diff-IP diff-EX diff-ER dist-1 dist-2 dist-3
Vicuna-7B 0.1857 0.2405 0.9509 0.3681 0.9745 0.2651 0.7995 0.9783 0.9936
ChatGLM3-6B 0.1632 0.2699 0.9768 0.344 0.9993 0.2629 0.7854 0.9776 0.9939
PEGS 0.1723 0.4458 0.9724 0.3113 0.8721 0.244 0.8397 0.9888 0.9976
Table 1: Results of quality of text generate in PEGS and baseline.
5.1 Experimental Settings
Implementation Details
At each training stage, the AdamW optimizer is adopted, and the cosine annealing scheduler is used to adjust the learning rate with the initial learning rate of 1e-4. The details of pre-training have been introduced in Section 4. Our model is fine-tuned on the entire StickerConv. It is noteworthy that only 3% of the model parameters are updated, and it takes about 6 hours on a single A800 GPU.
Baseline
We compare with the following baselines: (1) Vicuna-7B: An open-source chatbot trained by fine-tuning LLaMA on user-shared conversations collected from ShareGPT Chiang et al. (2023). We use the Vicuna-delta-v0 version to ensure consistency with the model base used during training. (2) ChatGLM3-6B: An open-source, bilingual supported large-scale dialogue model Zeng et al. (2023).
5.2 Evaluation Metrics
5.2.1 General Metrics
This category focuses on assessing the quality of the model’s textual output.
Diversity
We adopt Distinct-n (Dist-1, 2, 3) Li et al. (2016) as quantitative indicators of diversity, measuring the variety and creativity of the model’s outputs.
Empathy Performance
To evaluate the model’s capacity for empathy, we adopt the EPITOME Sharma et al. (2020a) method, comprising three sub-metrics: IP (interpretations), EX (explorations) and ER (emotional reactions).
Empathy Difference
DIFF-EPITOME Lee et al. (2022) measures the difference scores of IP, EX, ER between the golden response and predicted response.
5.2.2 LLM-based Metrics
LLM has been widely proven to be capable of acting as a grader similar to humans and can provide reasonable scores based on the designed prompts. These metrics not only focues on textual outputs but also include sticker outputs, forming a comprehensive multimodal evaluation system.
Empathy
Utilizing ChatGPT-3.5-turbo as a scorer, based on the context, we assess empathy scores for the model’s textual and sticker responses, denoted as Emp-txt and Emp-sticker, respectively. To mitigate randomness and bias, we conduct five independent scorings for each response and calculate the average.
Consistency
Again using ChatGPT-3.5-turbo as a scorer, based on the context, we assign consistency scores for textual and sticker responses, marked as Con-txt and Con-sticker, employing the same scoring method.
Ranking
Employing ChatGPT-3.5-turbo as a result ranker, we compare responses of different models to the same context. The ranker organizes responses based on quality, empathy, and consistency, averaging across all possible ranking combinations to ensure fairness and objectivity in assessment.
5.2.3 Multimodal Metrics
CLIP-Score
We utilize CLIP (Radford et al., 2021) as multimodal encoder, evaluate the consistency of model’s multimodal responses.
MM-Relevance
MM-Relevance Feng et al. (2023) is a metric designed to measure the relevance of multimodal responses. Traditional mm-relevance metrics, however, face limitations, particularly in cases where there is a lack of alignment between model responses and golden responses. To address these challenges, we propose Inter-MM-Rel (Interactive multimodal Relevance). This innovative approach calculates and averages the similarity scores between each mode of information, without the constraint of alignment, allowing for a more comprehensive and accurate assessment of multimodal response relevance.
5.3 Result Analysis
From Table 1, PEGS shows remarkable performance in several metrics, particularly in EX (Explorations), DIFF-EPITOME and diversity measuresm indicating a superior ability in generating diverse and exploratory responses. However, PEGS’s lower scores in IP, ER could be due to two reasons: (1) The general metrics only evaluate the text output and underestimate the actual empathy ability of the PEGS. (2) In multimodal empathy dialogue task, stickers play a significant role in conveying emotional reactions and explanations, while text handles more explorative tasks.
Models Rank Emp-txt Emp-sticker Con-txt Con-sticker
Vicuna 2.12 3.6768 4.3221
ChatGLM3 1.99 3.6908 4.3405
PEGS 1.88 3.7354 4.2534 4.2739 4.1155
Table 2: Results of LLM-Scorer.
Table 2 reveals PEGS’s strong performance in ranking, empathy text (Emp-txt), empathy sticker, and consistency of sticker responses (Con-sticker). However, its lower consistency in text (Con-txt) might be attributed to its emphasis on empathetic stickers, which could compromise some textual context consistency. This is in line with the additional information suggesting that the empathetic effectiveness of stickers might require some trade-offs in textual context consistency.
Models MM-Rel Inter-MM-Rel CLIP-Score
PEGS 69.41 59.06 21.03
Table 3: Results of multimodal evaluate.
Table 3 showcases the multimodal capabilities of PEGS. PEGS performs well in multimodal relevance (MM-Rel and Inter-MM-Rel) and consistency (CLIP-Score). This suggests that PEGS effectively integrates different modalities in its responses, maintaining relevance and consistency across them.
6 Conclusion
We explored the concept of multimodal empathetic response and created the StickerConv, the first dataset specifically designed for multimodal empathetic dialogue. We developed the Agent for StickerConv, a sophisticated LLM-based multi-agent system capable of simulating human-like interactions using stickers, thereby creating multimodal empathetic responses. Building upon the StickerConv, we developed PErceive and Generate Stickers (PEGS), an advanced multimodal dialogue model. This model is adept at both perceiving and generating stickers, effectively enhancing the communication experience. Furthermore, we established a comprehensive set of evaluation metrics for multimodal empathy tasks based on LLM. We are confident this work will be a valuable asset in advancing research in the field of multimodal empathetic dialogue systems.
Appendix A Appendix
A.1 Dataset Statistics
Table 4 offers a statistical breakdown of the dataset, segmented into training, validation, and test splits. It provides a quantitative overview, detailing the number of instances in each subset, the unique stickers, turn counts, and token counts.
split Number Unique Sticker Turn $\text{Token}_{\text{vicuna}}$ $\text{Token}_{\text{GPT}}$
train 10,785 4,798 59,424 10,681,108 9,070,221
validation 1,000 880 5,496 997,569 845,759
test 1,146 1,439 6,128 659,695 773,473
Table 4: The statistics of StickerConv.
Figure 5 showcases the emotion distribution within user profiles in the Agent for StickerConv. The chart illustrates the percentage share of each emotion, offering insight into the prevalence of affective expressions within the dataset.
Figure 5: Emotion distribution of user profile in Agent for StickerConv.
Figure 6, a word cloud, visualizes the 200 most prevalent emotion-related words extracted from the StickerConv. The prominence of each term in the cloud is indicative of its frequency within the dataset. This figure serves to underscore the linguistic diversity and the emotional range encapsulated in the dataset.
Figure 6: The 200 most popular emotion-related words in StickerConv.
Appendix B Case Study
We conduct a series of case studies to provide a more straightforward illustration of the performance of our model instruction-tuned by the StickerConv in multimodal empatheric conversations. These cases demonstrate the ability of PEGS to empathize with users and perceive the stickers provided by users, along with generating multimodal responses with vivid stickers.
As shown in Figure 7, in the positive emotion condition, PEGS will reward users with praise and support. As shown in Figure 8, two conversations in a negative scenario are displayed. PEGS as a conciliator can comfort users and guide their emotions towards positivity.
Appendix C Prompts
All {format_instructions} in the prompt utilize the response of langchain [^2] https://www.langchain.com/’s response_schema function implements format control. For detailed formatting information, please refer to our project’s open-source repository: https://github.com/ZhangYiqun018/StickerConv.
C.1 Sticker Process
Figure 9 shows the prompt of sticker process. For each sticker, we subject it to a process of filtering, annotation, and knowledge extraction, ultimately yielding three pieces of information: description, emotion and recommendation.
C.2 Sticker Agent Chat
Figure 10 shows the chat prompt, the summary encompasses the persona of the user or system, the current status, and the core traits generated based on memory. Figure 11, 12, 13 show the process of used sticker.
C.3 LLM-Based Scorer
Figure 14 illustrates the template for the preference rank scorer. This scorer integrates the outputs from all baseline models as well as PEGS, constituting the {response_list}. It then proceeds to rank these responses and provide reasoning for the assigned rankings.
Figures 15 and 16 depict the templates for the consistency scorer. The selection of the appropriate template depends on the modality of the response generated by the model. Specifically, Figure 15 is utilized for scoring purely textual responses. In contrast, when the response is multimodal, the scoring is conducted using the template shown in Figure 16.
Similarly, Figures 17 and 18 present the templates for the empathy scorer. The methodology for processing responses of different modalities aligns with the approach described for the consistency scorer.
C.4 Joint Learning
Figure 19 shows the instruction prompts for joint learning. We employed an instruction template similar to Vicuna_delta_v0. For every query or response that uses a sticker, we append an <IMG> identifier at the end.
Figure 7: Examples of conversations with positive emotions. Users can chat with multimodal content (text and stickers) and will receive positive multimodal responses.
Figure 8: Examples of conversations with negative emotions. Our model can empathize with users who are suffering from negative emotions, such as sadness (left) or anger (right) . Users will be comforted and guided to positive sentiments.
Figure 9: Prompt Templates for Process Sticker.
Figure 10: Prompt Template for Chat Prompt.
Figure 11: Prompt Template for Sticker Intent.
Figure 12: Prompt Template for Sticker Query.
Figure 13: Prompt Template for Sticker Select.
Figure 14: Prompt Template for Ranking score.
Figure 15: Prompt Template for Consistency Scorer (text only).
Figure 16: Prompt Template for Consistency Scorer (multimodal).
Figure 17: Prompt Template for Empathy Scorer (text only).
Figure 18: Prompt Template for Empathy Scorer (multimodal).
Figure 19: Prompt Template for Joint Learning.